# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v1

@@ Abstract
Artificial Intelligence increasingly pursues large, complex models that perform many tasks within increasingly realistic domains. How, if at all, should these developments in AI influence cognitive science? We argue that progress in AI offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes. We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms (and models that accommodate them) may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize. We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding. We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition—together with a reductive understanding of the processes and principles by which they do so.
@@ 1 Introduction
Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition and behavior (McClelland,, 2009). Cognitive models can range from simplified models of a high-level process (e.g. Frank and Goodman,, 2012; Cohen et al.,, 1990), to task-performing models of a particular domain (Newell,, 1973), or beyond (Newell,, 1994). Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001), and cognitive architectures (e.g. Ritter et al.,, 2019; Laird,, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles. Ultimately, these approaches could aspire to build towards ”unified theories of cognition” that can explain and predict natural behavior across many tasks and domains (Newell,, 1994). However, we are not there yet. In recent years there have been substantial debates about the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g. Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Wise et al.,, 2023; Cisek and Green,, 2024). Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
@@
The path we propose is motivated by recent practical and conceptual developments in AI. In the last ten years AI has progressed substantially in building models that can perform a wide range of tasks in various domains. We now have vision models that can classify and segment real-world images in objects without external supervision (He et al.,, 2020; Caron et al.,, 2021), we have reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and we have language models which can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022). Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al.,, 2014; Schrimpf et al.,, 2021). These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023). Yet they have also led to debate over how we test our models and interpret the results (Bowers et al.,, 2023; DiCarlo et al.,, 2023; Dentella et al.,, 2024). Thus, it remains an open question if and how the progress in AI can contribute to addressing the challenges of cognitive science.
@@
Figure 1: Naturalistic computational cognitive science: the what, why, and the how. The first section (§2) provides an overview of “naturalistic computational cognitive science”. Afterwards (§34), we motivate the utility of naturalistic experimental paradigms and learning-based approached for developing a more complete understanding of cognition. The rest of the paper (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories. (Figures reproduced from Saxe et al.,, 2019; Geirhos et al.,, 2018; Doshi and Konkle,, 2023.)
@@
In this work, we attempt to weave a thread of arguments that synthesizes these literatures, spanning from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility. We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve deeper understanding of the full range of natural intelligence. We present an overview of this framework and our arguments in Figure 1.
@@
Our paper proceeds as follows. First, in §2, we unpack what we mean by “naturalistic” computational cognitive science, by providing a pragmatic outline of the features that we believe deserve more emphasis in our experimental paradigms and models. In particular, we argue that researchers should expand their task paradigms to to more closely approximate the breadth of the settings in which their theoretical constructs would be expected to generalize by both adding relevant variables that may interact with those in question, and by incorporating more variability within existing parameters. Likewise, researchers should expand their models to more faithfully approximate the computational task that the natural system is solving.
@@
In the next two sections, we motivate these goals, by arguing that increasing naturalism is necessary for developing generalizable understanding. First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can lead to different behavior, engage brain systems differently, and expose expose computational challenges that engage mechanisms differently. As an example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes. Other works that we review have found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020) and more are changed when naturalistic elements are introduced. The works that we review in this section thus suggest that experiments that span a broader range of naturalistic settings are essential for drawing correct inferences about the system as a whole. Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can perform a wide range of tasks, and yield qualitatively different patterns of generalization than models trained in simplified settings. In turn, these differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena. Together, these two sections of our paper argue that pursuing naturalistic experiments, and models trained on them, will be necessary for achieving complete understanding of cognition.
@@
In the remainder of the paper, we turn to the practice of naturalistic computational cognitive science, and its contributions to building generalizable cognitive theories. First, in §5, we review key lessons computational cognitive scientists can adopt from AI in developing generalizable models that can operate in increasingly naturalistic data conditions: the benefits of frictionlessly reproducing prior research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions. Finally, in §6, we outline how one can leverage these potentially complicated and opaque models to contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021). We close in §7 by discussing the broader context and implications of naturalistic computational cognitive science.
@@ 2 What is “naturalistic” computational cognitive science?
Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict human behavior across increasingly naturalistic stimuli and tasks. This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions. As demonstrated in machine learning (§5), this strategy has enabled models that can work with real-world data, albeit imperfectly.
@@
    Changing existing experiment parameters to better cover the scope of natural distributions. For example, in an object classification task, we might expand the category set to include the broad range of objects people encounter in daily their life.
@@
Our goal is to encourage cognitive scientists to embrace experimental paradigms and models that capture a broader spectrum of the variability in inputs, interactions, and tasks present in the natural environment of humans or animals—including, for humans, the rich cultural constructs that form so much of our modern experience. That is, we identify a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize. Correspondingly, we identify a model as more naturalistic if it is capable of generalizing to make predictions across a broader range of contexts. Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments (see §6.1).
Generally, identifying the parameters for increasing naturalism in a particular experiment is not well defined or obvious—it is task- and theory-specific. To illustrate “naturalistic” more explicitly, we will therefore give some examples of dimensions along which naturalism can be increased.
@@
  • Incorporating multiple paradigms across which a hypothesis would be expected to hold, rather than testing on a single task; for example testing multi-step RL tasks as well as bandit tasks.
  • Incorporating a broader set of stimuli, for example augmenting synthetic images with real images, or images with videos.
  • Having task stimuli generated by many varying latent factors, not just the ostensible variables of theoretical interest.
@@
  • Including the presence of other social agents that continue to learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
  • A continually changing environment (Abel et al.,, 2024)
@@
  • Architectures that can operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g. low-dimensional or discrete inputs that contain only task-relevant features).
  • Architectures with naturalistic action spaces—for example, modeling low-level motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
@@
  • Prospective learning algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
Of course, it is easy to say that increasing the naturalism of our experiments and models would be useful, but how can we actually achieve it? A key goal for the later sections of our paper (§5-6) will be to give practical perspectives on how to achieve these goals, and how to develop theories that generalize across a broader range of naturalistic behavior. We briefly sketch these perspectives here.
How do we scale up to naturalistic experiments? We believe that recent developments in computer science and engineering make increasing naturalism in our experimental paradigms much more accessible to researchers, as we can use natural language to programmatically generate stimuli such as 3D scenes (Zhou et al.,, 2024) and videos of objects (Villegas et al.,, 2022), and leverage haptic feedback in virtual reality to simulate textures (Lu et al., 2022b, ). While the specific parameters will vary by domain, the core principle remains: gradually increasing ecological validity while maintaining experimental control.
How do we glean understanding from potentially opaque learning based models? (§6) We believe that task performing models offer powerful new tools for building cognitive theories. In a sense, these tools amount to treating the cognitive model as a proxy object for study—but one that admits much more opportunity for experimentation and understanding. For example, we can use “neuroscience”-like causal methods to probe a model’s computations by intervening on them (e.g. Geiger et al.,, 2021). We can also interpret its behavior as an adaptation to the naturalistic properties of its training data (Chan et al.,, 2022; Prystawski et al.,, 2024, e.g.)—as a kind of rational analysis (Anderson,, 1990; Lewis et al.,, 2014) of behavior and computations being a rational response to an environment. When applying these approaches to computational models, we have the ability to explicitly test our assumptions in ways we cannot with human subjects (e.g., how would the system behave if it were trained on only ungrammatical sentences). We can also attempt to distill our insights further, into analytic abstractions that illustrate the minimal instantiation of a phenomenon. By combining these routes, we can build rigorous links from an opaque task-performing model to conceptual understanding of cognition.
In summary naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over the scope of naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of an person’s environment, and (2) develop cognitive theories that unite these models with reductive understanding. In the coming sections, we will first motivate more deeply how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to the questions of how these goals can be achieved.
@@ 3 The benefits of naturalistic experimental settings
In this section, we outline the benefits of naturalistic experimental settings for cognitive science. Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby change the scientific inferences we draw. We review examples where naturalistic experimental settings have changed our understanding of a system, and have helped to disentangle competing models that simpler settings could not.
As an orienting conceptual example, imagine that we are doctors studying heart function, and we chose to only study it while the participants are lying down and resting. We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study the processes like how respiratory cycles affect heart rate. However, it offers this control and reliability precisely by removing the many factors of natural variation that interact with heart rate, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc. Studying heart function across a broader range of naturalistic settings would be necessary to understand why the system is the way it is, and how all the physiological (and psychological) processes involved interact. For example, if we experimentally manipulated breathing, but did so across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine how breathing affects heart function, and how that effect depends on the state of the overall system. That is, by studying phenomena across a range of naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.

### v2

@@ Abstract
How can cognitive science build generalizable theories that span the full scope of natural situations and behaviors? We argue that progress in Artificial Intelligence (AI) offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes. We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms, and models that accommodate them, may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize. First, we review cases from cognitive science and neuroscience where naturalistic paradigms elicit distinct behaviors or engage different processes. We then discuss recent progress in AI that shows that learning from naturalistic data yields qualitatively different patterns of behavior and generalization, and discuss how these findings impact the conclusions we draw from cognitive modeling, and can help yield new hypotheses for the roots of cognitive and neural phenomena. We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding. We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition—together with a reductive understanding of the processes and principles by which they do so.
Short Abstract
How should cognitive science pursue generalizable theories? We argue that progress in Artificial Intelligence (AI) offers opportunities to embrace naturalistic experiments and computational models that can accommodate them. We first review literature suggesting that building generalizable theories may require a broader range of naturalistic experimental paradigms. We then suggest that progress in AI enables incorporating naturalistic paradigms and models without losing experimental control or theoretically-grounded understanding. We argue that we are entering a new era in which cognitive science can build computational models that solve the real problems of natural cognition, and derive a reductive understanding of those solutions.
Keywords: Artificial intelligence, benchmarks, cognitive science, engineering, generalizability, naturalistic, task-performing models, theory
@@ 1 Introduction
Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition  (McClelland,, 2009). Cognitive models can range from simplified models of a high-level process (e.g. Frank and Goodman,, 2012; Cohen et al.,, 1990), to task-performing models of a particular domain (Newell,, 1973), or beyond (Newell,, 1994). Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001; Tenenbaum et al.,, 2011), cognitive architectures (e.g. Ritter et al.,, 2019; Laird,, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles. However, despite progress, there is still substantial debate on the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g. Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Nastase et al.,, 2020; Wise et al.,, 2023; Cisek and Green,, 2024). Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
@@
The path we propose is motivated by recent practical and conceptual developments in AI. In the last ten years, AI has progressed substantially in building models that can perform a wide range of tasks in various domains. We now have vision models that can classify and segment real-world images without external supervision (He et al.,, 2020; Caron et al.,, 2021), reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and language models that can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022). Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al.,, 2014; Schrimpf et al.,, 2021). These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ; Kanwisher et al.,, 2023), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023). Yet they have also led to debate over how we test our models and interpret the results (Bowers et al.,, 2023; DiCarlo et al.,, 2023; Dentella et al.,, 2024). Thus, it remains an open question if and how the progress in AI can contribute to addressing the challenges of cognitive science.
@@
Figure 1: Naturalistic computational cognitive science: the what, why, and the how. The first section (§2) provides an overview of “naturalistic computational cognitive science”. In (§34), we motivate using naturalistic experimental paradigms and learning-based approached for cognitive research. The remaining sections (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories. (Figures reproduced from Saxe et al.,, 2019; Geirhos et al.,, 2018; Doshi and Konkle,, 2023.)
@@
In this work, we attempt to weave a thread of arguments that synthesizes these literatures, spanning from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility. We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve a deeper understanding of the full range of natural intelligence. We present an overview of this framework and our arguments in Figure 1.
@@
Our paper proceeds as follows. First, in §2, we unpack what we mean by “naturalistic” computational cognitive science, by providing a pragmatic outline of the features that we believe deserve more emphasis in our experimental paradigms and models. We argue that researchers should expand their task paradigms to more closely approximate the breadth of the settings in which their theoretical constructs should be expected to generalize—first, by adding relevant variables that may interact with the theoretical constructs in question, and second by incorporating more variability within existing parameters. Likewise, researchers should expand their models to more faithfully approximate the computational task that the natural system is solving.
@@
In the next two sections, we motivate these goals by arguing that increasing naturalism is necessary for developing generalizable understanding. First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can yield different behavior, and alter how tasks engage computational mechanisms or neural systems. For example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes. Other work has found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020; Hasson et al.,, 2015) and more are changed when naturalistic elements are introduced. These works suggest that experiments that span a broader range of naturalistic settings can be essential for drawing correct inferences about the system as a whole. Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can yield qualitatively different patterns of generalization than models trained in simplified settings. These differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena. Together, these two sections of our paper argue that pursuing naturalistic experiments, and models that accommodate them, may be necessary for achieving a complete understanding of cognition.
@@
In the remainder of the paper, we turn to the practice of naturalistic computational cognitive science, and its contributions to building generalizable cognitive theories. First, in §5, we outline key lessons from AI in developing generalizable task-performing models for increasingly naturalistic data conditions: the benefits of being able to frictionlessly reproducing research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions. Finally, in §6, we outline how complicated experiments and opaque models can contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021). We close in §7 by discussing the broader context and implications of naturalistic computational cognitive science.
1.1 What this article is not
We first address some potential misconceptions we anticipate about our arguments.
First, we do not intend to argue that simple experimental paradigms are useless. Rather, we argue that incorporating more naturalistic settings alongside simpler ones is essential to developing more generalizable cognitive theories.
We also do not intend to promote a particular class of models, aside from our emphasis on learning. We believe that our arguments are compatible with many learning-based cognitive models, e.g. Bayesian cognitive models (Lee and Wagenmakers,, 2013), reinforcement learning (Collins,, 2019), predictive representations (Carvalho et al.,, 2024), program synthesis models (Rule et al.,, 2024), etc.
Finally, we do not advocate for black-box foundation models trained directly to reproduce experimental data as the goal of naturalistic computational cognitive science. While this can be a useful tool (e.g. for experimental design), we advocate for theory-driven development of predictive models that capture human behavior as an important tool in the quest to develop and validate generalizable cognitive theories.
In summary, this paper is not intended to reject the current experimental, modeling, and theoretical approaches of cognitive science. Rather, we advocate for extending from simpler experiments to more naturalistic ones. We then advocate for building task-performing models that can predict human behavior across these paradigms, and linking these models to simpler reductive models and theories of that behavior. We argue that this approach will enable building more generalizable theories that explicitly link cognitive capacities to the full scope of natural behavior to which those capacities contribute.
@@ 2 What is “naturalistic” computational cognitive science?
Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict and understand human behavior across increasingly naturalistic settings. This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions. In machine learning (§5), analogous strategies have enabled models that can work with real-world data, albeit imperfectly.
@@
    Expanding existing experiment parameters to better cover the scope of natural distributions. For example, in an object classification task, we might expand the category set to include the many objects people encounter in daily life.
@@
We describe a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize. We describe a model as more naturalistic if it is capable of generalizing its predictions across a broader range of contexts. Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments (see §6.1).
Generally, identifying the parameters for increasing naturalism in a particular experiment is not trivial—it is task- and theory-specific. To illustrate “naturalistic” more explicitly, we therefore offer some examples of dimensions along which naturalism can be increased.
@@
  • Incorporating multiple paradigms across which a hypothesis would be expected to hold, rather than relying on a single task; for example testing multi-step RL tasks as well as bandit tasks.
  • Incorporating broader stimulus sets, for example augmenting synthetic images with real images, or images with videos.
  • Having task stimuli generated by many varying latent factors, not just the variables of ostensible theoretical interest.
@@
  • Incorporating other social agents that learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
  • A continually changing environment (Abel et al.,, 2024).
@@
  • Architectures that operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g. low-dimensional or discrete inputs containing only task-relevant features).
  • Architectures with naturalistic action spaces—for example, modeling motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
@@
  • Prospective algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
In summary, naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of the environment, and (2) develop cognitive theories that unite these models with reductive understanding. In the coming sections, we will first motivate how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to how these goals can be achieved.
@@ 3 The benefits of naturalistic experimental settings
In this section, we outline the benefits of naturalistic experimental settings for cognitive science. Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby alter the scientific inferences that we draw. We review examples where naturalistic experimental settings have changed our understanding of a system, or have helped to disentangle competing models that simpler settings could not.
As an orienting conceptual example, imagine that we are doctors studying heart function, and we chose to only study it while the participants are lying down and resting. We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study processes like how respiration affects heart rate. However, this control and reliability comes precisely from removing the many factors of natural variation that affect the heart, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc. Studying heart function across a broader range of naturalistic settings would be necessary to understand the system in full, and how the physiological (and psychological) processes involved interact. For example, if we experimentally manipulated breathing across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine its effect on heart function, and how that effect varies across contexts. By studying phenomena across varied naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.